OpenStudy (anonymous):

Suppose the area of a square is 169 ft2. Find the perimeter.

OpenStudy (saifoo.khan):

\[\sqrt{169}=13\]

OpenStudy (saifoo.khan):

4(13) = answer.

OpenStudy (lalaly):

area=side^2 169=s^2 s=13 perimeter = 4s = 4*13

OpenStudy (lalaly):

=52

OpenStudy (saifoo.khan):

fts

OpenStudy (anonymous):

sq rt of 169 is 13 so 13*4 =52 52ft is the perimeter